# Two-Player Zero-Sum Games with Bandit Feedback

### v3

@@ Appendix C Proof of Theorem 2
According to the algorithm, $4\sigma\sqrt{e/T}$ is a critical value for $\hat{\Delta}_{t}$ since it takes this value in the round $\frac{1}{2}\log_{2}\frac{T}{e}$. In this way, let define a threshold parameter as $\lambda\geq 4\sigma\sqrt{e/T}$, which gives us two subsets of actions pairs such that $S_{A_{1}}=\{(i,j)\in S_{A}:\Delta_{ij}>\lambda\}$ and $S_{A_{2}}=\{(i,j)\in S_{A}:0<\Delta_{ij}\leq\lambda\}$. Since $\Delta_{ij}^{*}\leq\Delta_{ij}$, it ensures that the bounds are preserved and it makes sense to use these action pair sets in the analysis. Thus, we can write the Nash regret as
@@
$\varepsilon_{t_{ij}}=\sqrt{\frac{4\sigma^{2}}{k_{t_{ij}}}\ln{\Big(\frac{\hat{\Delta}_{t_{ij}}^{2}T}{16\sigma^{2}}\Big)}}\leq\frac{\hat{\Delta}_{t_{ij}}}{2}=\hat{\Delta}_{{t_{ij}}+1}<\frac{\Delta_{ij}}{4}$ (54)
where the exploration time $k_{t}$ is defined in (21), which satisfies $k_{t_{ij}}\geq\frac{16\sigma^{2}}{\hat{\Delta}_{t_{ij}}^{2}}\ln{\Big(\frac{\hat{\Delta}_{t}^{2}T}{16\sigma^{2}}\Big)}$.
@@
$\leq\exp{\Bigg(-\frac{k_{t}(\Delta_{ij}^{\text{max}}-\varepsilon_{t})^{2}}{4\sigma^{2}}\Bigg)}$ (60)
@@
$\leq\exp{\Bigg(-\frac{k_{t}(\Delta_{ij}^{\text{min}}-\varepsilon_{t})^{2}}{4\sigma^{2}}\Bigg)}$ (64)
@@
$\mathbb{P}(E_{1}\cap E_{2})\leq\min\{\mathbb{P}(E_{1}),\mathbb{P}(E_{2})\}$ $\leq\min\Bigg\{\exp{\Bigg(-\frac{k_{t}(\Delta_{ij}^{\text{max}}-\varepsilon_{t})^{2}}{4\sigma^{2}}\Bigg)},\exp{\Bigg(-\frac{k_{t}(\Delta_{ij}^{\text{min}}-\varepsilon_{t})^{2}}{4\sigma^{2}}\Bigg)}\Bigg\}$ (65)
$=\exp{\Big(-\frac{k_{t}\max\{(\Delta_{ij}^{\text{max}}-\varepsilon_{t})^{2},(\Delta_{ij}^{\text{min}}-\varepsilon_{t})^{2}\}}{4\sigma^{2}}\Big)}$ (66)
$\leq\exp{\Big(-\frac{k_{t}((\Delta_{ij}^{\text{max}}-\varepsilon_{t})^{2}+(\Delta_{ij}^{\text{min}}-\varepsilon_{t})^{2})}{8\sigma^{2}}\Big)}$ (67)
$\leq\exp{\Big(-\frac{k_{t}(\Delta_{ij}-2\varepsilon_{t})^{2}}{16\sigma^{2}}\Big)}$ (68)
$<\exp{\Big(-\frac{k_{t}\varepsilon_{t}^{2}}{4\sigma^{2}}\Big)}$ (69)
$=\exp{\Bigg(-\frac{k_{t}\frac{4\sigma^{2}}{k_{t}}\ln{\Big(\frac{\hat{\Delta}_{t}^{2}T}{16\sigma^{2}}\Big)}}{4\sigma^{2}}\Bigg)}$ (70)
@@
$k_{t_{ij}}=\Big\lceil\frac{16\sigma^{2}}{\hat{\Delta}_{t_{ij}}^{2}}\ln{(\frac{\hat{\Delta}_{t_{ij}}^{2}T}{16\sigma^{2}})}\Big\rceil\leq\Big\lceil\frac{256\sigma^{2}}{\Delta_{ij}^{2}}\ln{\Big(\frac{\Delta_{ij}^{2}T}{256\sigma^{2}}\Big)}\Big\rceil.$ (73)
@@
$\sum_{(i,j)\in S_{A_{1}}}\Delta_{ij}^{*}\Big\lceil\frac{256\sigma^{2}}{\Delta_{ij}^{2}}\ln{\Big(\frac{\Delta_{ij}^{2}T}{256\sigma^{2}}\Big)}\Big\rceil$ $\leq\sum_{(i,j)\in S_{A_{1}}}\Delta_{ij}^{*}\Big(1+\frac{256\sigma^{2}}{\Delta_{ij}^{2}}\ln{\Big(\frac{\Delta_{ij}^{2}T}{256\sigma^{2}}\Big)}\Big)$ (74)
$=\sum_{(i,j)\in S_{A_{1}}}\Delta_{ij}^{*}+\Delta_{ij}^{*}\frac{256\sigma^{2}}{\Delta_{ij}^{2}}\ln{\Big(\frac{\Delta_{ij}^{2}T}{256\sigma^{2}}\Big)}$ (75)
@@
We note that $\Delta_{ij}^{*}\leq\Delta_{ij}$, which ensures that the regret does not grow excessively, and this relationship allows us to analyze the regret in terms of the defined action sets. As a result, if we sum these regret contributions as mentioned, we can conclude the proof. $\Box$
@@ Appendix D Proof of Theorem 4
$\leq\min\Bigg\{\exp{\Bigg(-\frac{(\Delta_{ij}^{\text{max}}-\varepsilon_{t})^{2}}{2(\frac{\sigma^{2}}{k_{i^{\prime}j,t}}+\frac{\sigma^{2}}{k_{ij,t}})}\Bigg)},\exp{\Bigg(-\frac{(\Delta_{ij}^{\text{min}}-\varepsilon_{t})^{2}}{2(\frac{\sigma^{2}}{k_{ij^{\prime},t}}+\frac{\sigma^{2}}{k_{ij,t}})}\Bigg)}\Bigg\}$ (80)
$\leq\min\Bigg\{\exp{\Bigg(-\frac{(\Delta_{ij}^{\text{max}}-\varepsilon_{t})^{2}}{2\Big(\frac{\sigma^{2}}{(k_{t}+k_{t}\hat{P}_{i^{\prime}j,t})}+\frac{\sigma^{2}}{(k_{t}+k_{t}\hat{P}_{ij,t})}\Big)}\Bigg)},\exp{\Bigg(-\frac{(\Delta_{ij}^{\text{min}}-\varepsilon_{t})^{2}}{2\Big(\frac{\sigma^{2}}{(k_{t}+k_{t}\hat{P}_{ij^{\prime},t})}+\frac{\sigma^{2}}{(k_{t}+k_{t}\hat{P}_{ij,t})}\Big)}\Bigg)}\Bigg\}$ (81)
$\leq\min\Bigg\{\exp{\Bigg(-\frac{(\Delta_{ij}^{\text{max}}-\varepsilon_{t})^{2}}{2\sigma^{2}\Big(\frac{(\hat{P}_{i^{\prime}j,t}+\hat{P}_{ij,t}+2)}{k_{t}(1+\hat{P}_{i^{\prime}j,t})(1+\hat{P}_{ij,t})}\Big)}\Bigg)},\exp{\Bigg(-\frac{(\Delta_{ij}^{\text{min}}-\varepsilon_{t})^{2}}{2\sigma^{2}\Big(\frac{(\hat{P}_{ij^{\prime},t}+\hat{P}_{ij,t}+2)}{k_{t}(1+\hat{P}_{ij^{\prime},t})(1+\hat{P}_{ij,t})}\Big)}\Bigg)}\Bigg\}$ (82)
$\leq\min\Bigg\{\exp{\Bigg(-\frac{k_{t}(\Delta_{ij}^{\text{max}}-\varepsilon_{t})^{2}}{4\sigma^{2}}\Bigg)},\exp{\Bigg(-\frac{k_{t}(\Delta_{ij}^{\text{min}}-\varepsilon_{t})^{2}}{4\sigma^{2}}\Bigg)}\Bigg\}$ (83)
@@
$k_{t_{ij}}=\Big\lceil\frac{16\sigma^{2}}{\hat{\Delta}_{t_{ij}}^{2}}\ln{(\frac{\hat{\Delta}_{t_{ij}}^{2}T}{16\sigma^{2}})}\Big\rceil+\Big\lceil\frac{16\sigma^{2}\hat{P}_{ij,t_{ij}}}{\hat{\Delta}_{t_{ij}}^{2}}\ln{(\frac{\hat{\Delta}_{t_{ij}}^{2}T}{16\sigma^{2}})}\Big\rceil\leq 2\ \Big\lceil\frac{256\sigma^{2}}{\Delta_{ij}^{2}}\ln{\Big(\frac{\Delta_{ij}^{2}T}{256\sigma^{2}}\Big)}\Big\rceil$ (85)
@@
$\sum_{(i,j)\in S_{A_{1}}}\Delta_{ij}^{*}\ 2\ \Big\lceil\frac{256\sigma^{2}}{\Delta_{ij}^{2}}\ln{\Big(\frac{\Delta_{ij}^{2}T}{256\sigma^{2}}\Big)}\Big\rceil$ $\leq\sum_{(i,j)\in S_{A_{1}}}\Delta_{ij}^{*}+\Delta_{ij}^{*}\frac{512\sigma^{2}}{\Delta_{ij}^{2}}\ln{\Big(\frac{\Delta_{ij}^{2}T}{256\sigma^{2}}\Big)}$ (86)
@@
$R_{T}\leq\sum_{(i,j)\in S_{A_{1}}}\Delta_{ij}+\frac{768\sigma^{2}}{\Delta_{ij}}+\frac{512\sigma^{2}}{\Delta_{ij}}\ln{\Big(\frac{\Delta_{ij}^{2}T}{256\sigma^{2}}\Big)}\Big)+\sum_{(i,j)\in S_{A_{2}}}\Big(\frac{512\sigma^{2}}{\lambda}+\lambda T\Big).$
Since we have $\Delta_{ij}^{*}\leq\Delta_{ij}$, it concludes the proof.

### v4

@@ Appendix C Proof of Theorem 2
According to the algorithm, $4\sigma\sqrt{e/T}$ is a critical value for $\hat{\Delta}_{t}$ since it takes this value in the round $\frac{1}{2}\log_{2}\frac{T}{e}$. In this way, let define a threshold parameter as $\lambda\geq 4\sigma\sqrt{e/T}$, which gives us two subsets of actions pairs such that $S_{A_{1}}=\{(i,j)\in S_{A}:\Delta_{ij}>\lambda\}$ and $S_{A_{2}}=\{(i,j)\in S_{A}:0<\Delta_{ij}\leq\lambda\}$. Since $\Delta_{ij}^{*}\leq\Delta_{ij}$ from Lemma 1, it ensures that the bounds are preserved and it makes sense to use these action pair sets in the analysis. Thus, we can write the Nash regret as
@@
$\varepsilon_{t_{ij}}=\sqrt{\frac{4\sigma^{2}}{k_{t_{ij}}}\ln{\left(\frac{\hat{\Delta}_{t_{ij}}^{2}T}{16\sigma^{2}}\right)}}\leq\frac{\hat{\Delta}_{t_{ij}}}{2}=\hat{\Delta}_{{t_{ij}}+1}<\frac{\Delta_{ij}}{4}$ (50)
where the exploration time $k_{t}$ is defined in (18), which satisfies $k_{t_{ij}}\geq\frac{16\sigma^{2}}{\hat{\Delta}_{t_{ij}}^{2}}\ln{\left(\frac{\hat{\Delta}_{t}^{2}T}{16\sigma^{2}}\right)}$.
@@
$\leq\exp{\left(-\frac{k_{t}(\Delta_{ij}^{\text{max}}-\varepsilon_{t})^{2}}{4\sigma^{2}}\right)}$ (56)
@@
$\leq\exp{\left(-\frac{k_{t}(\Delta_{ij}^{\text{min}}-\varepsilon_{t})^{2}}{4\sigma^{2}}\right)}$ (60)
@@
$\mathbb{P}(E_{1}\cap E_{2})\leq\min\{\mathbb{P}(E_{1}),\mathbb{P}(E_{2})\}$ $\leq\min\left\{\exp{\left(-\frac{k_{t}(\Delta_{ij}^{\text{max}}-\varepsilon_{t})^{2}}{4\sigma^{2}}\right)},\exp{\left(-\frac{k_{t}(\Delta_{ij}^{\text{min}}-\varepsilon_{t})^{2}}{4\sigma^{2}}\right)}\right\}$ (61)
$=\exp{\left(-\frac{k_{t}\max\{(\Delta_{ij}^{\text{max}}-\varepsilon_{t})^{2},(\Delta_{ij}^{\text{min}}-\varepsilon_{t})^{2}\}}{4\sigma^{2}}\right)}$ (62)
$\leq\exp{\left(-\frac{k_{t}((\Delta_{ij}^{\text{max}}-\varepsilon_{t})^{2}+(\Delta_{ij}^{\text{min}}-\varepsilon_{t})^{2})}{8\sigma^{2}}\right)}$ (63)
$\leq\exp{\left(-\frac{k_{t}(\Delta_{ij}-2\varepsilon_{t})^{2}}{16\sigma^{2}}\right)}$ (64)
$<\exp{\left(-\frac{k_{t}\varepsilon_{t}^{2}}{4\sigma^{2}}\right)}$ (65)
$=\exp{\left(-\frac{k_{t}\frac{4\sigma^{2}}{k_{t}}\ln{\left(\frac{\hat{\Delta}_{t}^{2}T}{16\sigma^{2}}\right)}}{4\sigma^{2}}\right)}$ (66)
@@
$k_{t_{ij}}=\left\lceil\frac{16\sigma^{2}}{\hat{\Delta}_{t_{ij}}^{2}}\ln{(\frac{\hat{\Delta}_{t_{ij}}^{2}T}{16\sigma^{2}})}\right\rceil\leq\left\lceil\frac{256\sigma^{2}}{\Delta_{ij}^{2}}\ln{\left(\frac{\Delta_{ij}^{2}T}{256\sigma^{2}}\right)}\right\rceil.$ (69)
@@
$\sum_{(i,j)\in S_{A_{1}}}\Delta_{ij}^{*}\left\lceil\frac{256\sigma^{2}}{\Delta_{ij}^{2}}\ln{\left(\frac{\Delta_{ij}^{2}T}{256\sigma^{2}}\right)}\right\rceil$ $\leq\sum_{(i,j)\in S_{A_{1}}}\Delta_{ij}^{*}\left(1+\frac{256\sigma^{2}}{\Delta_{ij}^{2}}\ln{\left(\frac{\Delta_{ij}^{2}T}{256\sigma^{2}}\right)}\right)$ (70)
$=\sum_{(i,j)\in S_{A_{1}}}\Delta_{ij}^{*}+\Delta_{ij}^{*}\frac{256\sigma^{2}}{\Delta_{ij}^{2}}\ln{\left(\frac{\Delta_{ij}^{2}T}{256\sigma^{2}}\right)}$ (71)
@@
Lemma 1 ensures that the regret does not grow excessively, and this relationship allows us to analyze the regret in terms of the defined action sets. As a result, if we sum these regret contributions as mentioned, we can conclude the proof. $\Box$
@@ Appendix D Proof of Theorem 4
$\leq\min\left\{\exp{\left(-\frac{(\Delta_{ij}^{\text{max}}-\varepsilon_{t})^{2}}{2(\frac{\sigma^{2}}{k_{i^{\prime}j,t}}+\frac{\sigma^{2}}{k_{ij,t}})}\right)},\exp{\left(-\frac{(\Delta_{ij}^{\text{min}}-\varepsilon_{t})^{2}}{2(\frac{\sigma^{2}}{k_{ij^{\prime},t}}+\frac{\sigma^{2}}{k_{ij,t}})}\right)}\right\}$ (76)
$\leq\min\left\{\exp{\left(-\frac{(\Delta_{ij}^{\text{max}}-\varepsilon_{t})^{2}}{2\left(\frac{\sigma^{2}}{(k_{t}+k_{t}\hat{P}_{i^{\prime}j,t})}+\frac{\sigma^{2}}{(k_{t}+k_{t}\hat{P}_{ij,t})}\right)}\right)},\exp{\left(-\frac{(\Delta_{ij}^{\text{min}}-\varepsilon_{t})^{2}}{2\left(\frac{\sigma^{2}}{(k_{t}+k_{t}\hat{P}_{ij^{\prime},t})}+\frac{\sigma^{2}}{(k_{t}+k_{t}\hat{P}_{ij,t})}\right)}\right)}\right\}$ (77)
$\leq\min\left\{\exp{\left(-\frac{(\Delta_{ij}^{\text{max}}-\varepsilon_{t})^{2}}{2\sigma^{2}\left(\frac{(\hat{P}_{i^{\prime}j,t}+\hat{P}_{ij,t}+2)}{k_{t}(1+\hat{P}_{i^{\prime}j,t})(1+\hat{P}_{ij,t})}\right)}\right)},\exp{\left(-\frac{(\Delta_{ij}^{\text{min}}-\varepsilon_{t})^{2}}{2\sigma^{2}\left(\frac{(\hat{P}_{ij^{\prime},t}+\hat{P}_{ij,t}+2)}{k_{t}(1+\hat{P}_{ij^{\prime},t})(1+\hat{P}_{ij,t})}\right)}\right)}\right\}$ (78)
$\leq\min\left\{\exp{\left(-\frac{k_{t}(\Delta_{ij}^{\text{max}}-\varepsilon_{t})^{2}}{4\sigma^{2}}\right)},\exp{\left(-\frac{k_{t}(\Delta_{ij}^{\text{min}}-\varepsilon_{t})^{2}}{4\sigma^{2}}\right)}\right\}$ (79)
@@
$k_{t_{ij}}=\left\lceil\frac{16\sigma^{2}}{\hat{\Delta}_{t_{ij}}^{2}}\ln{\left(\frac{\hat{\Delta}_{t_{ij}}^{2}T}{16\sigma^{2}}\right)}\right\rceil+\left\lceil\frac{16\sigma^{2}\hat{P}_{ij,t_{ij}}}{\hat{\Delta}_{t_{ij}}^{2}}\ln{\left(\frac{\hat{\Delta}_{t_{ij}}^{2}T}{16\sigma^{2}}\right)}\right\rceil\leq 2\ \left\lceil\frac{256\sigma^{2}}{\Delta_{ij}^{2}}\ln{\left(\frac{\Delta_{ij}^{2}T}{256\sigma^{2}}\right)}\right\rceil$ (81)
@@
$\sum_{(i,j)\in S_{A_{1}}}\Delta_{ij}^{*}\ 2\ \left\lceil\frac{256\sigma^{2}}{\Delta_{ij}^{2}}\ln{\left(\frac{\Delta_{ij}^{2}T}{256\sigma^{2}}\right)}\right\rceil$ $\leq\sum_{(i,j)\in S_{A_{1}}}\Delta_{ij}^{*}+\Delta_{ij}^{*}\frac{512\sigma^{2}}{\Delta_{ij}^{2}}\ln{\left(\frac{\Delta_{ij}^{2}T}{256\sigma^{2}}\right)}$ (82)
@@
$R_{T}\leq\sum_{(i,j)\in S_{A_{1}}}\Delta_{ij}+\frac{768\sigma^{2}}{\Delta_{ij}}+\frac{512\sigma^{2}}{\Delta_{ij}}\ln{\left(\frac{\Delta_{ij}^{2}T}{256\sigma^{2}}\right)}+\sum_{(i,j)\in S_{A_{2}}}\left(\frac{512\sigma^{2}}{\lambda}+\lambda T\right).$
Since we have $\Delta_{ij}^{*}\leq\Delta_{ij}$ from Lemma 1, it concludes the proof.